Frances fired a projectile from the origin (at y=0 m) on top of a cliff. The initial velocity components are Viy=230 m/s and Vix=96m/s. The time to reach the ground is 36s, and the horizontal distance traveled is 3.5 km. The y coordinate of the highest point reached by the projectile is:

a) 470m b) 2,700m c) 12,300m d) 940m e) 90,160m
Answer #1

first of all see that you need a vector because you are working on 2D. So \theta =tan(\frac{y}{x}) then Vo=(x^{2}+y^{2})^{\frac{1}{2}}

so in the X plane you have

Vx=96m/s

a=0

x=vo.cos(\theta ).t

and in the plane Y

Vy=230m/s

ay=-g

y=vo.sin(\theta ).t-\frac{1}{2}.g.t

then if you calculate the y max you get.

ymax=\frac{vo^{2}.sen^{2}(\theta )}{g} where g=9.81 \frac{m}{s^{2}} find and that's set